A passenger plane made a trip to Las Vegas and back. On the trip there it flew 432 mph and on the return trip it went 480 mph. How long did the trip there take if the return trip took nine hours?
Answer by geetha_rama(94)

Onward trip speed = 432 mph
Return speed = 480 mph
Time taken in return trip = 9 hrs
distance = speed * time
= 480 * 9 = 4320 miles
Time taken for the onward trip = distance/onward speed = 4320/432 = 10 hrs